OpenStudy (artfis):

Ok so I got this question wrong. What is the way to solve this system of equations? {2(y - 6) = 3y + 4(x-5) {2(4x + 3) - 5 = 2(1 - y) + 5x

OpenStudy (anonymous):

whenever you simplified the equations what did you get

OpenStudy (artfis):

I distributed the numbers that were near the parentheses first.

OpenStudy (anonymous):

after simplification i got first equation: y + 4x = 8 second equation : 3x+2y = 1 then y = 8 - 4x so putting the first into the second 3x + 2(8 - 4x) = 1 3x + 16 - 8x = 1 -5x = -15 x = 3 y = 8 - 4(x) y = 8 - 4(3) y = 8 - 12 y = -4

OpenStudy (anonymous):

first equation: 2(y - 6) = 3y + 4(x-5) 2y - 12 = 3y + 4x - 20 -12 + 20 = 4x + 3y - 2y 8 = 4x + y or y + 4x = 8 second equation 2(4x + 3) - 5 = 2(1 - y) + 5x 8x + 6 - 5 = 2 - 2y + 5x 8x - 5x + 2y = 2 + 5 - 6 3x + 2y = 1

OpenStudy (artfis):

........Oh ok you make it seem so clear. I don't know how I wsas mixing it up. Must be too much stress. Thank you so much @Davidjohn

OpenStudy (anonymous):

no problem, lots of practice. It's easy to mix up a number when there are so many